OpenStudy (anonymous):

In 4^x+3=8^x-1, what is the value of x?

OpenStudy (anonymous):

@dan815 how do you go about doing this?

OpenStudy (kirbykirby):

\[\ln4^x+3=8^x-1 \]?

OpenStudy (anonymous):

I think it is \[ 4^{x+3}=8^{x-1}\]

OpenStudy (kirbykirby):

Ok that is much easier to do then lol

OpenStudy (anonymous):

yes then it is..

OpenStudy (anonymous):

No, the first one. @kirbykirby

OpenStudy (anonymous):

It's actually: \[4^{x} + 3 = 8^{x} - 1\]

OpenStudy (anonymous):

I bet kirby is thinking Ln, and you are thinking in

OpenStudy (anonymous):

Without the ln.

OpenStudy (anonymous):

Yeah.

OpenStudy (kirbykirby):

ooh ok lol

OpenStudy (experimentx):

if you are looking for integer solution ... then factorize 8^x - 1 ... and use divisibility, it will give you one solution x=1

OpenStudy (experimentx):

for x=2 and higher ... show that 8^x-1 > 4^x +2 hence no solution exist. for negative x, use similar argument.

OpenStudy (anonymous):

Oh, thank you.